OpenStudy (anonymous):

Determine the values of x where the tangent line is horizontal for the function: f(x) = \frac{1}{(x^2 - 25)( x - 7)}. The value(s) of x where the tangent line to the graph of the function is horizontal is(are) .

OpenStudy (ash2326):

Find f'(x) and equate it to 0, to find the value of x

OpenStudy (amistre64):

latex only parses with the delimiters capping it

OpenStudy (anonymous):

\[f(x) = \frac{1}{(x^2 - 25)( x - 7)}.\]

OpenStudy (ash2326):

for which tangent is horizontal or slope is zero

OpenStudy (amistre64):

i almost thought of this as a max/min and was gonna say f'(x)=0 wasnt sufficient; but, as is it is correct to find horizontal tangents ;)

OpenStudy (anonymous):

i know i have to find the derivative of f(x) which is the same as the tangent line

OpenStudy (anonymous):

and then from the drivative solve for x

OpenStudy (amistre64):

since a quotient rule still has to carry out the product rule; id just rewrite for product

OpenStudy (anonymous):

but i tried this and didnt get the correct answer

OpenStudy (anonymous):

what a pain. derivative is \[f'(x)=\frac{-3x^2+14x+25}{(x^2-25)^2(x-7)^2}\]so you have to set \[-3x^2+14x+25=0\] and use quadratic formula

OpenStudy (anonymous):

so now i dont know what to do

OpenStudy (anonymous):

so i got x = -1.37850958 and 6.04517624

OpenStudy (amistre64):

ewww, id keep it in exact form

OpenStudy (anonymous):

should i round these numbers?

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

so what do i do after this

OpenStudy (amistre64):

you move on to the next problem. a good rule of thumb is: when you get to the end, stop.

OpenStudy (anonymous):

but dont i have to find a y value too? how can i ?

OpenStudy (amistre64):

"Determine the values of x where..."; so dont go creating any extra work for ourself unless you are really really excited

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

thanks... i only needed the x values and math doesnt make me excited ...its just so difficult

OpenStudy (amistre64):

;) good luck